# Testing Stochastic Block Models Based on Maximum Sampling Entry-Wise Deviations

### v1

@@ 1 Introduction
Due to the rapid development and application of information technology, immense troves of network data have been created across various fields, such as finance and business (Katona et al. 2011 and Di Maggio et al. 2019), health care (Valente 2010; Lawyer 2015), engineering (Agarwal et al. 2013), machine learning (Musumeci et al. 2018) and social science (Tabassum et al. 2018). To make full use of network data, it is critical to understand their generation mechanisms. To this end, the stochastic block model (SBM) has been widely studied and used since the pioneering work of Holland et al. (1983); see, e.g., Snijders and Nowicki (1997), Nowicki and Snijders (2001), Bickel and Chen (2009), Rohe et al. (2011), Choi et al. (2012), Jin (2015), Zhang and Zhou (2016), and Hu et al. (2021). A good review paper on SBM can be found in Lee and Wilkinson (2019).
@@
The aim of this paper is to develop a new test statistic that allows for testing the number of communities $K$ and the mapping vector $g=(g(1),\cdots,g(n))^{\top}$, without regard to the sparsity of the network or the divergence of $K$. To obtain this test, we not only allow $K$ to be divergent, but also simply require the condition, $n\min_{i,j}Q_{g(i)g(j)}\rightarrow\infty$. This is a relaxed version of the Condition (ii) mentioned earlier. The procedures of the proposed method are as follows. We first adopt the approach of Hu et al. (2021) to calculate the entry-wise deviation $\widehat{\rho}_{iv}$ for each node $i$ in community $v\in\{1,\cdots,K\}$, which is constructed based on the average of the standardized summation of $a_{ij}$ for all nodes $j\not=i$ within the community $v$. Subsequently, for any fixed community $v\in\{1,\cdots,K\}$, we randomly draw $B$ samples from $n$ nodes with replacement and then calculate $\widehat{\gamma}_{\cdot v}=\sum_{i=1}^{B}\widehat{\rho}_{iv}/\sqrt{B}$. Repeating this sampling procedure $M$ times for each of the $K$ communities, we obtain the $M$ values of $\widehat{\gamma}_{\cdot v}$, and denote them $\widehat{\gamma}_{mv}$ for $m=1,\cdots,M$. Finally, the new test statistic, $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\gamma}_{mv}|$, is constructed. By the central limit theorem, we can demonstrate that $\widehat{\gamma}_{mv}$ converges to the standard normal distribution $N(0,1)$ plus an additional term of the order $O_{p}(1/\sqrt{B\log n})$. By setting $B$ to be divergent, the additional term with the order $O_{p}(1/\sqrt{B\log n})$ becomes negligible. As a result, we expect that $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\gamma}_{mv}|$ converges to an extreme value distribution under certain conditions regardless of dense networks or sparse networks.
@@
Although implementing our proposed test is not complicated, establishing its theoretical properties is challenging. This is because the $\widehat{\gamma}_{mv}$’s are correlated due to the replacement samples from $\widehat{\rho}_{iv}$. As a result, $\max_{m,v}|\widehat{\gamma}_{mv}|$ is no longer the maximum of a sequence of independent variables, and its asymptotic distribution is more difficult to demonstrate than that of Hu et al. (2021). In the paper, we overcome challenges and show that $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\gamma}_{mv}|$ converges to the Type-I extreme value distribution. In addition, according to the aforementioned explanations, our approach comprises Hu et al. (2021)’s method and the sampling with replacement procedure. Accordingly, our proposed method is versatile for both dense and sparse networks. In contrast, Hu et al. (2021)’s method is tailored for dense networks. Consequently, our proposed test exhibits greater robustness in practical application. Since the test statistic’s rate of convergence to the extreme distribution may be slow, we further consider a finite-sample bootstrap correction by generating bootstrap samples to improve the finite sample performance. Finally, we introduce an augmented test statistic to increase the power and extend the proposed test to the degree-corrected SBM.
@@ 2.3 Maximum sampling entry-wise deviation test statistic
$\widehat{\gamma}_{mv}=\frac{\widehat{\rho}_{m_{1}v}+\widehat{\rho}_{m_{2}v}+\cdots+\widehat{\rho}_{m_{B}v}}{\sqrt{B}}.$ (2.3)
Repeating this procedure $M$ times for each community $v$, we obtain the sequence of $\widehat{\gamma}_{mv}$’s for $m=1,\cdots,M$. Under the null hypotheses in (I) and (II), one expects all $\widehat{\rho}_{iv}$ values to be small as noted by Hu et al. (2021). Accordingly, $\widehat{\gamma}_{mv}$ should also be small, which motivates us to propose the maximum sampling entry-wise deviation based on the maximum value of $\widehat{\gamma}_{mv}$’s across $m$ and $v$,
$F_{n}(K_{0},g_{0})\overset{\Delta}{=}\max_{1\leq m\leq M,1\leq v\leq K_{0}}|\widehat{\gamma}_{mv}|.$ (2.4)
By setting $B=1$ and choosing $M$ to be sufficiently large such that all the $\widehat{\rho}_{iv}$’s can be selected, $F_{n}(K_{0},g_{0})$ is then identical to $L_{n}(K_{0},g_{0})$.
It is worth noting that the above sampling process can alleviate the negative impact of network sparsity by letting $B$ be divergent in $n$. Considering a sparse network with $c_{\min}n^{-1}\log n\leq\min_{i,j}Q_{g(i)g(j)}\leq\max_{i,j}Q_{g(i)g(j)}\leq c_{\max}n^{-1}\log n$ as discussed in Section 2.2, $\widehat{\rho}_{iv}$ is a summation of $c_{\rho}\log n$ random terms. Accordingly, (2.3) can be expressed as a summation of infinite $c_{\rho}B\log n$ random terms as $B\to\infty$ along with $n$. By the central limit theorem, $\widehat{\gamma}_{mv}$ can be expressed as the standard normal distribution $N(0,1)$ plus a term $O_{p}(1/\sqrt{B\log n})$. By Lemma 1 given in the supplementary material, we can show that $F_{n}(K_{0},g_{0})=O_{p}\big{\{}\sqrt{\log(MK_{0})}\big{\}}\big{\{}O_{p}(1)+O_{p}(1/\sqrt{B\log n})\big{\}}=O_{p}\big{\{}\sqrt{\log(MK_{0})}\big{\}}+o_{p}(1)$ under Conditions (C2) and (C3) below. Thus, the term $O_{p}(1/\sqrt{B\log n})$ is negligible in deriving the asymptotic distribution of $F_{n}(K_{0},g_{0})$. As a result, we expect that $F_{n}(K_{0},g_{0})$ follows an extreme value distribution under certain conditions, and it is applicable to both dense and sparse networks. To rigorously derive the theoretical properties of $F_{n}(K_{0},g_{0})$, we introduce the following technical conditions.
@@
Condition (C1) requires that the probability matrix $Q$ is identifiable with entries all bounded in $(0,1)$. This condition is similar to, but weaker than, that of Lei (2016), Wang and Bickel (2017), and Hu et al. (2021). Note that if all $Q_{g(i)g(j)}$s have the same order, then $Q_{g(i)g(j)}$ shares the same order with the network density $D_{n}$. Accordingly, this condition immediately implies that the network density $D_{n}$ satisfies $nD_{n}\rightarrow\infty$ and allows the network density to approach zero more slowly than $n^{-1}$ as $n\to\infty$. In addition, if $\widehat{Q}^{g}_{g(i)g(j)}=0$, then the term $\big{(}a_{ij}-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}\big{)}/\big{(}\sqrt{\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}(1-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}})}\big{)}$ in the summation of (2.2) is removed, and our analysis remains valid. For the sake of simplicity, we omit the case which $Q_{g(i)g(j)}=0$ in this paper. Condition (C2) is a mild condition that requires the size of the smallest community to be at least proportional to $n/K$. Furthermore, Condition (C2) imposes an upper bound on the largest community size, which was also considered by Zhang and Zhou (2016), Gao et al. (2018), and Hu et al. (2021). This condition is used to control the estimation bias between $\widehat{Q}^{g}_{g(i)g(j)}$ and its population version $Q_{g(i)g(j)}$, $\max_{1\leq i\leq n,1\leq v\leq K}\left|\frac{1}{\sqrt{|g^{-1}(v)/\{i\}|}}\sum_{j\in g^{-1}(v)/\{i\}}\frac{Q_{g(i)g(j)}-\widehat{Q}^{g}_{g(i)g(j)}}{\sqrt{Q_{g(i)g(j)}(1-Q_{g(i)g(j)})}}\right|.$ Specifically, the condition ensures that this estimation bias converges to zero in probability. Condition (C3) identifies the orders of $K$, $B$, and $M$. Intuitively, $B$ and $M$ cannot be too large; otherwise, the correlation between $\widehat{\gamma}_{mv}$’s may be large. Note that Condition (C3)-(ii) allows $K$ to grow more slowly along $n$ than the assumption that $K=o{(n/\log^{2}n)}$, which was proposed by Hu et al. (2021). In addition, Han et al. (2023) required $K$ to be a fixed value, although they introduced less stringent density conditions than those in Lei (2016), Wang and Bickel (2017), and Hu et al. (2021). Consequently, in comparison to Lei (2016), Wang and Bickel (2017), Hu et al. (2021), and Han et al. (2023), our method not only exhibits the adaptability to a broad spectrum of network densities, but also allows $K$ to be divergent. Based on the above three conditions, we next obtain the theoretical properties of $T_{n}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0}).$
@@
$\lim_{n\to\infty}P\Big{\{}T_{n}\leq x\Big{\}}=\exp\big{(}-\frac{1}{\sqrt{\pi}}e^{-x/2}\big{)},$ (2.5)
as $n\to\infty$, where the right hand side of (2.5) is the CDF of the Type-I extreme value distribution (i.e., Gumbel distribution) with $\mu=-2\log(\sqrt{\pi})$ and $\beta=2$.
Applying Theorem 1, we can test the null hypotheses in (I) and (II). To test the null hypothesis in (I), we replace $g_{0}$ with its consistent estimator $\widehat{g}$. This estimator can be obtained regardless the network being dense, $n\min_{i,j}Q_{g(i)g(j)}/\log n\rightarrow\infty$, (see, e.g., Daudin et al. 2008; Lei and Rinaldo 2015; Lei and Zhu 2017; Gao et al. 2017 and Gao et al. 2018) or sparse, $n\max_{i,j}Q_{g(i)g(j)}=O(\log n)$, (see, e.g., Zhang and Zhou 2016; Jing et al. 2022). The resulting test statistic is $T_{n,1}=F_{n}^{2}(K_{0},\widehat{g})-2\log(MK_{0})+\log\log(MK_{0}).$ We can demonstrate theoretically that $T_{n,1}$ has the same asymptotic null distribution as $T_{n}$ under proper conditions; see Corollary S.1 and its proof in Section S.7 of the supplementary material. Consequently, we reject the null hypothesis, $H_{0k}:K=K_{0}$, if $T_{n,1}>q_{\alpha}$, where $q_{\alpha}=-2\log\{-\sqrt{\pi}\log(1-\alpha)\}$ is the $(1-\alpha)$-th quantile of the Type-I extreme distribution with the distribution function $\exp\big{(}-e^{-x/2}/\sqrt{\pi}\big{)}$.
As for testing the null hypothesis in (II), we can directly calculate the test statistic: $T_{n,2}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0}).$ This is because $g_{0}$ can yield its corresponding $K_{0}$. Thus, we can reject the null hypothesis $H_{0g}:g=g_{0}$ when $T_{n,2}>q_{\alpha}$ at the given significance level $\alpha$.
For ease of understanding and practical applications, we present Algorithm 1 below with five steps to calculate the test statistic $T_{n,2}$ for testing (II). To obtain $T_{n,1}$ for testing (I), one can replace $g_{0}$ in $T_{n,2}$ with its consistent estimator $\widehat{g}$ calculated under $K=K_{0}$.
@@
Output: Test statistic $T_{n,2}$ and the decision in regard to the null hypothesis $H_{0g}$ being rejected or not.
@@
4At the $m$-th realization for $m=1,\cdots M$, randomly draw $B$ samples from $\{\widehat{\rho}_{1v},\cdots,\widehat{\rho}_{nv}\}$ obtained in Step 2 for each $1\leq v\leq K_{0}$, and calculate $\widehat{\gamma}_{mv}$ using (2.3).
5Find the maximum value of $\widehat{\gamma}_{mv}$ for $1\leq m\leq M$ and $1\leq v\leq K_{0}$, and obtain $F_{n}(K_{0},g_{0})$ using (2.4).
Calculate $T_{n,2}=F_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0})$ and reject the null hypothesis $H_{0g}:g=g_{0}$ if $T_{n,2}>q_{\alpha}$.

### v2

# A Goodness-of-Fit Test for Sparse Networks

@@ 1 Introduction
Due to the rapid development and application of information technology, immense troves of network data have been created across various fields, such as finance and business (Katona et al. 2011; Di Maggio et al. 2019; Härdle et al. 2016), social science (Bramoullé et al. 2009; Jin et al. 2024; Andrikopoulos et al. 2016) engineering (Agarwal et al. 2013), machine learning (Musumeci et al. 2018). To make full use of network data, it is critical to understand their generation mechanisms. To this end, the stochastic block model (SBM) has been widely studied and used since the pioneering work of Holland et al. (1983); see, e.g., Snijders and Nowicki (1997), Nowicki and Snijders (2001), Bickel and Chen (2009), Rohe et al. (2011), Choi et al. (2012), Jin (2015), Zhang and Zhou (2016), Hu et al. (2021) and Jochmans (2024). A good review paper on SBM can be found in Lee and Wilkinson (2019).
@@
The aim of this paper is to develop a new test statistic that allows for testing the number of communities $K$ and the mapping vector $g=(g(1),\cdots,g(n))^{\top}$, without regard to the sparsity of the network or the divergence of $K$. To obtain this test, we not only allow $K$ to be divergent, but also simply require the condition, $n\min_{i,j}Q_{g(i)g(j)}\rightarrow\infty$. This is a relaxed version of the Condition (ii) mentioned earlier. The procedures of the proposed method are as follows. We first adopt the approach of Hu et al. (2021) to calculate the entry-wise deviation $\widehat{\rho}_{iv}$ for each node $i$ in community $v\in\{1,\cdots,K\}$, which is constructed based on the average of the standardized summation of $a_{ij}$ for all nodes $j\not=i$ within the community $v$. Subsequently, for any fixed community $v\in\{1,\cdots,K\}$, we randomly draw $B$ samples from $n$ nodes with replacement and then calculate $\widehat{\psi}_{\cdot v}=\sum_{i=1}^{B}\widehat{\rho}_{iv}/\sqrt{B}$. Repeating this sampling procedure $M$ times for each of the $K$ communities, we obtain the $M$ values of $\widehat{\psi}_{\cdot v}$, and denote them $\widehat{\psi}_{mv}$ for $m=1,\cdots,M$. Finally, the new test statistic, $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\psi}_{mv}|$, is constructed. By the central limit theorem, we can demonstrate that $\widehat{\psi}_{mv}$ converges to the standard normal distribution $N(0,1)$ plus an additional term of the order $O_{p}(1/\sqrt{B\log n})$. By setting $B$ to be divergent, the additional term with the order $O_{p}(1/\sqrt{B\log n})$ becomes negligible. As a result, we expect that $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\psi}_{mv}|$ converges to an extreme value distribution under certain conditions regardless of dense networks or sparse networks.
@@
Although implementing our proposed test is not complicated, establishing its theoretical properties is challenging. This is because the $\widehat{\psi}_{mv}$’s are correlated due to the replacement samples from $\widehat{\rho}_{iv}$. As a result, $\max_{m,v}|\widehat{\psi}_{mv}|$ is no longer the maximum of a sequence of independent variables, and its asymptotic distribution is more difficult to demonstrate than that of Hu et al. (2021). In the paper, we overcome challenges and show that $\max_{1\leq m\leq M,1\leq v\leq K}|\widehat{\psi}_{mv}|$ converges to the Type-I extreme value distribution. In addition, according to the aforementioned explanations, our approach comprises Hu et al. (2021)’s method and the sampling with replacement procedure. Accordingly, our proposed method is versatile for both dense and sparse networks. In contrast, Hu et al. (2021)’s method is tailored for dense networks. Consequently, our proposed test exhibits greater robustness in practical application. Since the test statistic’s rate of convergence to the extreme distribution may be slow, we further consider a finite-sample bootstrap correction by generating bootstrap samples to improve the finite sample performance. Finally, we introduce an augmented test statistic to increase the power and extend the proposed test to the degree-corrected SBM.
@@ 2.2 Maximum entry-wise deviation
Remark 1.
It is worth noting that the entrywise deviation statistic $\widehat{\rho}_{iv}$, proposed by Hu et al. (2021), has also been extended to various network testing problems, such as two-sample testing; see, for example, Wu et al. (2024a), Wu et al. (2024) and Fu et al. (2025). These methods are tailored for dense networks (Fu et al., 2025; Wu et al., 2024a), or for sparse regimes satisfying $\min_{i,j}Q_{g(i)g(j)}\geq c_{\min}n^{-1}\log n$ for a constant $c_{\min}>0$ (Wu et al., 2024). However, it is important to note that these methods are not applicable to the testing problems considered in our work, as they focus on assessing whether two networks differ in their community structures. In these works, they assume that both networks are already generated from a SBM with a pre-specified number of communities $K$. Thus, these works are not capable of simultaneously testing the two hypotheses (I) and (II) considered in our paper.
@@ 2.3 Maximum sampling entry-wise deviation test statistic
$\widehat{\psi}_{mv}=\frac{1}{\sqrt{B}}\sum_{h=1}^{B}\widehat{\rho}_{m_{h}v}.$ (2.3)
Repeating this procedure $M$ times for each community $v$, we obtain the sequence of $\widehat{\psi}_{mv}$’s for $m=1,\cdots,M$. Under the null hypotheses in (I) and (II), one expects all $\widehat{\rho}_{iv}$ values to be small as noted by Hu et al. (2021). Accordingly, $\widehat{\psi}_{mv}$ should also be small, which motivates us to propose the maximum sampling entry-wise deviation based on the maximum value of $\widehat{\psi}_{mv}$’s across $m$ and $v$,
$\Gamma_{n}(K_{0},g_{0})\overset{\Delta}{=}\max_{1\leq m\leq M,1\leq v\leq K_{0}}|\widehat{\psi}_{mv}|.$ (2.4)
By setting $B=1$ and choosing $M$ to be sufficiently large such that all the $\widehat{\rho}_{iv}$’s can be selected, $\Gamma_{n}(K_{0},g_{0})$ is then identical to $L_{n}(K_{0},g_{0})$.
It is worth noting that the above sampling process can alleviate the negative impact of network sparsity by letting $B$ be divergent in $n$. Considering a sparse network with $c_{\min}n^{-1}\log n\leq\min_{i,j}Q_{g(i)g(j)}\leq\max_{i,j}Q_{g(i)g(j)}\leq c_{\max}n^{-1}\log n$ as discussed in Section 2.2, $\widehat{\rho}_{iv}$ is a summation of $c_{\rho}\log n$ random terms. Accordingly, (2.3) can be expressed as a summation of infinite $c_{\rho}B\log n$ random terms as $B\to\infty$ along with $n$. By the central limit theorem, $\widehat{\psi}_{mv}$ can be expressed as the standard normal distribution $N(0,1)$ plus a term $O_{p}(1/\sqrt{B\log n})$. By Lemma 1 given in the supplementary material, we can show that $\Gamma_{n}(K_{0},g_{0})=O_{p}\big{\{}\sqrt{\log(MK_{0})}\big{\}}\big{\{}O_{p}(1)+O_{p}(1/\sqrt{B\log n})\big{\}}=O_{p}\big{\{}\sqrt{\log(MK_{0})}\big{\}}+o_{p}(1)$ under Conditions (C2) and (C3) below. Thus, the term $O_{p}(1/\sqrt{B\log n})$ is negligible in deriving the asymptotic distribution of $\Gamma_{n}(K_{0},g_{0})$. As a result, we expect that $\Gamma_{n}(K_{0},g_{0})$ follows an extreme value distribution under certain conditions, and it is applicable to both dense and sparse networks. To rigorously derive the theoretical properties of $\Gamma_{n}(K_{0},g_{0})$, we introduce the following technical conditions.
@@
Condition (C1) requires that the probability matrix $Q$ is identifiable with entries all bounded in $(0,1)$. This condition is similar to, but weaker than, that of Lei (2016), Wang and Bickel (2017), and Hu et al. (2021). Note that if all $Q_{g(i)g(j)}$s have the same order, then $Q_{g(i)g(j)}$ shares the same order with the network density $D_{n}$. Accordingly, this condition immediately implies that the network density $D_{n}$ satisfies $nD_{n}\rightarrow\infty$ and allows the network density to approach zero more slowly than $n^{-1}$ as $n\to\infty$. In addition, if $\widehat{Q}^{g}_{g(i)g(j)}=0$, then the term $\big{(}a_{ij}-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}\big{)}/\big{(}\sqrt{\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}}(1-\widehat{Q}_{g_{0}(i)g_{0}(j)}^{g_{0}})}\big{)}$ in the summation of (2.2) is removed, and our analysis remains valid. For the sake of simplicity, we omit the case which $Q_{g(i)g(j)}=0$ in this paper. Condition (C2) is a mild condition that requires the size of the smallest community to be at least proportional to $n/K$. Furthermore, Condition (C2) imposes an upper bound on the largest community size, which was also considered by Zhang and Zhou (2016), Gao et al. (2018), and Hu et al. (2021). This condition is used to control the estimation bias between $\widehat{Q}^{g}_{g(i)g(j)}$ and its population version $Q_{g(i)g(j)}$, $\max_{1\leq i\leq n,1\leq v\leq K}\left|\frac{1}{\sqrt{|g^{-1}(v)/\{i\}|}}\sum_{j\in g^{-1}(v)/\{i\}}\frac{Q_{g(i)g(j)}-\widehat{Q}^{g}_{g(i)g(j)}}{\sqrt{Q_{g(i)g(j)}(1-Q_{g(i)g(j)})}}\right|.$ Specifically, the condition ensures that this estimation bias converges to zero in probability. Condition (C3) identifies the orders of $K$, $B$, and $M$. Intuitively, $B$ and $M$ cannot be too large; otherwise, the correlation between $\widehat{\psi}_{mv}$’s may be large. Note that Condition (C3)-(ii) allows $K$ to grow more slowly along $n$ than the assumption that $K=o{(n/\log^{2}n)}$, which was proposed by Hu et al. (2021). In addition, Han et al. (2023) required $K$ to be a fixed value, although they introduced less stringent density conditions than those in Lei (2016), Wang and Bickel (2017), and Hu et al. (2021). Consequently, in comparison to Lei (2016), Wang and Bickel (2017), Hu et al. (2021), and Han et al. (2023), our method not only exhibits the adaptability to a broad spectrum of network densities, but also allows $K$ to be divergent. Based on the above three conditions, we next obtain the theoretical properties of $\Theta_{n}=\Gamma_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0}).$
@@
$\lim_{n\to\infty}P\Big{\{}\Theta_{n}\leq x\Big{\}}=\exp\big{(}-\frac{1}{\sqrt{\pi}}e^{-x/2}\big{)},$ (2.5)
where the right hand side of (2.5) is the CDF of the Type-I extreme value distribution (i.e., Gumbel distribution) with $\mu=-2\log(\sqrt{\pi})$ and $\beta=2$.
Applying Theorem 1, we can test the null hypotheses in (I) and (II). To test the null hypothesis in (I), we replace $g_{0}$ with its consistent estimator $\widehat{g}$. This estimator can be obtained regardless the network being dense, $n\min_{i,j}Q_{g(i)g(j)}/\log n\rightarrow\infty$, (see, e.g., Daudin et al. 2008; Lei and Rinaldo 2015; Lei and Zhu 2017; Gao et al. 2017 and Gao et al. 2018) or sparse, $n\max_{i,j}Q_{g(i)g(j)}=O(\log n)$, (see, e.g., Zhang and Zhou 2016; Jing et al. 2022). The resulting test statistic is $\Theta_{n,1}=\Gamma_{n}^{2}(K_{0},\widehat{g})-2\log(MK_{0})+\log\log(MK_{0}).$ We can demonstrate theoretically that $\Theta_{n,1}$ has the same asymptotic null distribution as $\Theta_{n}$ under proper conditions; see Corollary S.1 and its proof in Section S.7 of the supplementary material. Consequently, we reject the null hypothesis, $H_{0k}:K=K_{0}$, if $\Theta_{n,1}>q_{\alpha}$, where $q_{\alpha}=-2\log\{-\sqrt{\pi}\log(1-\alpha)\}$ is the $(1-\alpha)$-th quantile of the Type-I extreme distribution with the distribution function $\exp\big{(}-e^{-x/2}/\sqrt{\pi}\big{)}$.
As for testing the null hypothesis in (II), we can directly calculate the test statistic: $\Theta_{n,2}=\Gamma_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0}).$ This is because $g_{0}$ can yield its corresponding $K_{0}$. Thus, we can reject the null hypothesis $H_{0g}:g=g_{0}$ when $\Theta_{n,2}>q_{\alpha}$ at the given significance level $\alpha$.
For ease of understanding and practical applications, we present Algorithm 1 below with five steps to calculate the test statistic $\Theta_{n,2}$ for testing (II). To obtain $\Theta_{n,1}$ for testing (I), one can replace $g_{0}$ in $\Theta_{n,2}$ with its consistent estimator $\widehat{g}$ calculated under $K=K_{0}$.
@@
Output: Test statistic $\Theta_{n,2}$ and the decision in regard to the null hypothesis $H_{0g}$ being rejected or not.
@@
4At the $m$-th realization for $m=1,\cdots M$, randomly draw $B$ samples from $\{\widehat{\rho}_{1v},\cdots,\widehat{\rho}_{nv}\}$ obtained in Step 2 for each $1\leq v\leq K_{0}$, and calculate $\widehat{\psi}_{mv}$ using (2.3).
5Find the maximum value of $\widehat{\psi}_{mv}$ for $1\leq m\leq M$ and $1\leq v\leq K_{0}$, and obtain $\Gamma_{n}(K_{0},g_{0})$ using (2.4).
Calculate $\Theta_{n,2}=\Gamma_{n}^{2}(K_{0},g_{0})-2\log(MK_{0})+\log\log(MK_{0})$ and reject the null hypothesis $H_{0g}:g=g_{0}$ if $\Theta_{n,2}>q_{\alpha}$.